# Relying on the Unreliable: The Impact of Language Models’ Reluctance to Express Uncertainty

### v1

@@ Abstract
As natural language becomes the default interface for human-AI interaction, there is a critical need for LMs to appropriately communicate uncertainties in downstream applications. In this work, we investigate how LMs incorporate confidence about their responses via natural language and how downstream users behave in response to LM-articulated uncertainties. We examine publicly deployed models and find that LMs are unable to express uncertainties when answering questions even when they produce incorrect responses. LMs can be explicitly prompted to express confidences, but tend to be overconfident, resulting in high error rates (on average 47%) among confident responses. We test the risks of LM overconfidence by running human experiments and show that users rely heavily on LM generations, whether or not they are marked by certainty. Lastly, we investigate the preference-annotated datasets used in RLHF alignment and find that humans have a bias against texts with uncertainty. Our work highlights a new set of safety harms facing human-LM interactions and proposes design recommendations and mitigating strategies moving forward.
@@
[…]
@@ 1 Introduction
A pivotal aspect of fostering reliable human-AI interactions lies in the apt communication of model uncertainties Cai et al. (2019); Kizilcec (2016); De-Arteaga et al. (2020), typically defined as the probability assigned to a model’s prediction. Recent work in language generation reflects a shift towards using natural language as a means to convey model confidences (e.g., “I’m fairly confident it’s”, "According to Wikipedia it’s" Mielke et al., 2022; Lin et al., 2022; Zhou et al., 2023). Such features are known in the linguistics literature as epistemic markers, which serve to convey a speaker’s stance and commitment, which support human communication and decision making Babrow et al. (1998); Brashers et al. (2000); Tseng and Zhang (2023). Since epistemic markers play an important role in human-human communication and decision making Budescu et al. (1988); Windschitl and Wells (1996); Druzdzel (1989), we hypothesize that the use of these markers by LMs will also have an impact on human-AI interactions.
@@
Our work begins with an examination of how LMs communicate uncertainties to end-users in realistic information seeking scenarios (§3). Specifically, we elicit responses from popular, publicly-deployed models including GPT, LLaMA-2, and Claude by prompting them to provide epistemic markers when answering multiple choice questions Brown et al. (2020); OpenAI (2023); Touvron et al. (2023); Anthropic (2022). Our analysis reveals that LMs are reluctant to share model uncertainties, despite errors in their generations. LMs can be explicitly prompt to use epistemic markers, but are more likely to generate expressions of certainty than uncertainty, despite an average 47% error rate among high confidence responses, (e.g., I’m confident the capital of Tanzania is Dar es Salaam. [Incorrect]).
We then investigate the behavioral responses of end-users to model-generated epistemic markers (§4). While linguists and psychologists have long focused on the interpretation of epistemic markers by humans Budescu et al. (1988); Windschitl and Wells (1996); Wallsten et al. (1986), the pragmatic implications of speakers of epistemic markers being AI systems, combined with dangers of over-reliance on AI Bussone et al. (2015); Jacobs et al. (2021); Bansal et al. (2021b); Buçinca et al. (2021) could drastically change their interpretation compared to human-spoken ones. We address this gap by conducting several user studies to measure how individuals interpret and respond to uncertainties as articulated by LMs in calibrated and miscalibrated settings (Figure 1). Our findings indicate that users are heavily reliant on LM generated expressions of high confidence (e.g., "I’m sure it’s…"), but are also surprisingly reliant on plain statements (e.g., "The answer is…"). Additionally, we find that even minor miscalibrations in how a model uses epistemic markers can lead to long-term harms in human performance.
Lastly, given our findings on model overconfidence and human reliability of LM generations, we pinpoint the origins of model overconfidence (§5). We investigate model artifacts such as base models, instruction-tuned models, reward models, and human feedback datasets to isolate the origins of model confidence. Our investigation identifies the process of reinforcement learning with human feedback (RLHF) as a key contributing factor and we uncover that human annotators are biased against expressions of uncertainty.
@@ 2 Epistemic Markers in Language Models
Linguists has extensively studied epistemic markers as ubiquitous linguistic features that signal the speaker’s commitment and stance. These markers broadly fall into two categories: weakeners—expressions of uncertainty, and strengtheners—expressions of certainty. Numerous other categories include: hedges Lakoff (1975), boosters Hyland (2005, 2014), evidentials Aikhenvald (2004), approximators Prince et al. (1982), and factive verbs Kiparsky and Kiparsky (1970).
In machine learning, work has focused on improving model calibration Jiang et al. (2021); Desai and Durrett (2020); Jagannatha and Yu (2020); Kamath et al. (2020); Kong et al. (2020) by calibrating the confidence value assigned by a model and model accuracy through a measure called ECE Naeini et al. (2015). Recent work has focused explicitly on how pretraining Hendrycks et al. (2019) and scaling Srivastava et al. (2022); Chen et al. (2022) impacts the calibration of language models. Most relevant to our work is Dhuliawala et al. (2023)’s studies on how humans interpret numerical confidences in calibrated and miscalibrated settings. A key challenge remains, as numeric confidence values are known to be challenging for users to interpret Miller (2019). Our work aims to fill a gap by providing a more comprehensive understanding of how humans interpret LM-generated epistemic markers.
We begin by eliciting open-ended generations, a departure from prior methods which prompts models to produce a predefined set of confidence expressions either numerically Kadavath et al. (2022); Tian et al. (2023); Liu et al. (2023); Xiong et al. (2023); Tanneru et al. (2023) or verbally (based on an ordinal scale) Lin et al. (2022); Mielke et al. (2022). This strategy enables a bottom-up approach towards understanding how LMs generate epistemic markers and more closely mimics how real-world deployed users might engage with chat models.
In addition to examining what LMs generate, we also reveal how users behave in response to these generations in both static and interactive settings. By focusing on how humans interpret these markers, we provide a framework for how to measure the harms of linguistic miscalibration from the lens of downstream users.
@@ 3.1 Methods
Our objective is to assess the potential harms and safety risks associated with widely used publicly deployed models like GPT, LLaMA-2, and Claude. [^1] Models were accessed June - November 2023. We pose a diverse set of questions from the Massive Multitask Language Understanding benchmark (MMLU Hendrycks et al., 2021), a dataset comprised of multiple-choice questions with four options each. These questions span 57 subjects that assess both language model knowledge and problem-solving skills. In this section, we design confidence eliciting prompts and measure if there are systematic trends in how LMs use strengtheners and weakeners.
@@
To ensure the generalizability of our results, we employ snowball sampling to generate a list of diverse prompts, gathering additional paraphrases of prompts from Amazon Mechanical Turk Workers and GPT-3.5 (details in A).
We prompted nine models (text-davinci-003, GPT-3.5-Turbo, GPT-4, LLaMA-2 7B, LLaMA-2 13B, LLaMA-2 70B, Claude-1, Claude-2, Claude- Instant-1) using 49 prompts on 284 questions, resulting in a total of 125,244 queries. [^2] Duplicated question in the development set of MMLU benchmark, resulting in 284 instead of 285 questions. Using a zero-shot prompting approach, we aim to simulate the interactions of end-users. We set the temperature to 0.3 OpenAI (2023); Anil et al. (2023), with no stop tokens, and limit the token generation length to a maximum of 400 tokens (Details in C).
@@
Eliciting and Classifying Epistemic Markers
The authors then qualitatively code for epistemic markers in generated responses, iteratively identifying the markers within the responses and categorizing them as strengtheners and weakeners. Using Regex pattern matching, we automatically identify epistemic markers in response generations. [^3] A future iteration could include training a classifier to identify codes at the trade-off of lowered transparency and interpretability. In contrast to previous analyses that primarily examined single-word lexical forms of epistemic markers like “anyway”, “should,” and “obviously” Islam et al. (2020), our analysis specifically focuses on phrase-levels of epistemic markers. Our qualitative coding yielded a total of 76 strengtheners and 105 weakeners. See tables 8 and 9 for the most commonly generated expressions.
3.2 Findings
@@
When using chain-of-thought instructions or explicit instructions to elicit epistemic markers, LMs can be encouraged to produce more epistemic markers. Resulting in 16% and 71% of generations incorporating epistemic markers respectively.
@@
Six out of nine models have a preference to generate significantly more strengtheners than weakeners. Our results indicate that an average of 20% of generations had strengtheners meanwhile only 14% had weakeners, and this bias is true among prompts eliciting for certainty, with or without CoT (Table 1). We see this trend emerge strongly among GPT and LLaMA-2 chat models, with the Claude-2 being more balanced in its generation of strengtheners and weakeners (Figure 3). Interestingly, we find that smaller models (LLaMA-2-7B and Claude-Instant-V1444Announced as a ”lighter” version of Claude. See: https://www.anthropic.com/index/introducing-claude) have a higher use of weakeners over strengtheners, contrasting with larger models.
@@
We find that across all generations, only 53% of generations with expressions of certainty are correct (random guessing accuracy being 25%). Although this accuracy rate is higher than accuracies among weakeners (32%), this is an alarmingly high rate of errors among strengtheners. Furthermore due to the high rate of generations of strengtheners, 17% of all incorrect answers include strengtheners.
@@ 3.3 Discussion
In the space of information seeking, the need to express uncertainties and knowledge limitations will likely continue to persist. The need for linguistic calibration is emerging as a new harm as LMs play a bigger role in human-LM collaborations and will likely persist even as model performance improves. Moving forward, we must examine how to mitigate the harms of model overconfidence and how to best build cognitive forcing designs, such as verbalized uncertainties, to discourage human overreliance of AI systems Buçinca et al. (2021).
@@ 4.1 Methods
For details on recruitment process and template selection see Appendix B.
@@ 4.2 Experimental Settings
Figure 4: Example of human experiments task for Setting 1.
Figure 5: Participant results in the calibrated, overconfident, and underconfident settings. We see lowered scores across the miscalibrated rounds but in the overconfident setting, lowered scored persist in the later calibrated rounds as well.
@@
We first use our task to evaluate how participants rely on LM generated epistemic markers (from §3). Participants are shown a set of trivia questions and the beginning of a response (e.g., "I think it’s…" Figure 4). Users are asked whether or not they’d like to rely on Marvin’s answer, and since the users do not see any answers, they are simply expressing their reliance of epistemic markers as generated by LMs. Participants are presented with strengtheners, weakeners, and plain statements, which are expressions free of epistemic markers like, “The answer is (A)”. We recruited 25 participants and each were shown 106 questions.
@@
In these experimental settings, we introduce a scoring system, also modified from Bansal et al. 2019a. The scoring set-up is designed such that the only way to have a positive score is to rely on Marvin correctly based on reading the epistemic markers (see Table 2). [^7] To avoid spammers, we filtered the bottom 20% participants based on their performance on this task.
@@ 4.3 Findings
When users are presented with weakeners, approximately 90% of users will choose to look up the answer themselves. When users are presented with expressions of strengtheners, nearly 90%, users will rely on Marvin’s response (See tables 6 and 7 for details). However, surprisingly, plain statements, which are expressions free of epistemic markers like, “The answer is (A)” or “(A)”, are relied on by users nearly 90% of the time. In other words, without communicating any epistemic markers, humans interpret this as a sign of model certainty.
@@
In the calibrated setting, our results illustrate that users are able to learn mental models of epistemic markers after approximately 20 rounds. In the early stages, users rely on strengtheners 94% of the time and on weakeners 7% of the time. After 20 rounds, most participants are able to nearly perfectly leverage Marvin’s epistemic markers with users learning to rely on strengtheners 99% of the time and weakeners nearly 1% of the time, averaging an accuracy of 97% across all rounds. The high performance in this calibrated setting also validates that participants are primarily relying on epistemic markers, to answer these questions. If they had been only relying on their own knowledge, it would be unlikely to see participants perform nearly perfectly on such challenging trivia questions.
@@
Participants in the overconfident setting are too reliant of strengtheners. In the first 20 rounds, users relied on 81% of strengtheners, when only 66% of the generations with strengtheners were correct, accruing on greater penalties than necessary. Because participants are unlikely to know the answer to the question, we see that participants mistakenly rely on incorrect, confident generations an average 73% of the time. The consequences of system miscalibration continued to negatively impact human performance, even after several rounds of calibrated model answers. Participants averaged 76% in the miscalibrated rounds and 86% on the calibrated rounds (Figure 5).
@@
An unexpected effect of the overconfident responses was that participants started to interpret weakeners differently. In the overconfident setting, strengtheners were used incorrectly but weakeners were used correctly (only occurring when the answer is wrong). However, we see participants respond to weakeners differently, with users relying on answers with weakeners nearly 10%, when none of the answers with weakeners were correct.
@@
In the underconfident settings, users are not reliant enough of weakeners — 33% of generations with weakeners were correct but users only relied on those generations 17% of the time. Participants in the underconfident setting averaged 66% in the miscalibrated rounds but 98% on the calibrated rounds, matching the performance in the calibrated setting. This is in contrast to the overconfident setting where users’ mental models formed correctly, even after the same number of calibrated rounds.

### v2

@@ Abstract
As natural language becomes the default interface for human-AI interaction, there is a need for LMs to appropriately communicate uncertainties in downstream applications. In this work, we investigate how LMs incorporate confidence in responses via natural language and how downstream users behave in response to LM-articulated uncertainties. We examine publicly deployed models and find that LMs are reluctant to express uncertainties when answering questions even when they produce incorrect responses. LMs can be explicitly prompted to express confidences, but tend to be overconfident, resulting in high error rates (an average of 47%) among confident responses. We test the risks of LM overconfidence by conducting human experiments and show that users rely heavily on LM generations, whether or not they are marked by certainty. Lastly, we investigate the preference-annotated datasets used in post training alignment and find that humans are biased against texts with uncertainty. Our work highlights new safety harms facing human-LM interactions and proposes design recommendations and mitigating strategies moving forward.
@@
Kaitlyn Zhou1,4   Jena D. Hwang4   Xiang Ren2,4   Maarten Sap3,4 1Stanford University, 2University of Southern California, 3Carnegie Mellon University, 4Allen Institute for AI katezhou@stanford.edu
@@ 1 Introduction
A pivotal aspect of fostering reliable human-AI interactions lies in the apt communication of model uncertainties Cai et al. (2019); Kizilcec (2016); De-Arteaga et al. (2020), typically defined as the probability assigned to a model’s prediction. Recent work in language generation reflects a shift towards using natural language as a means to convey model confidences (e.g., “I’m fairly confident it’s”, "According to Wikipedia it’s" Mielke et al., 2020; Lin et al., 2022; Zhou et al., 2023). Such features are known in the linguistics literature as epistemic markers, which serve to convey a speaker’s stance and commitment, thus supporting human communication and decision making Babrow et al. (1998); Brashers et al. (2000); Tseng and Zhang (2023). Since epistemic markers play an important role in human-human communication and decision making Budescu et al. (1988); Windschitl and Wells (1996); Druzdzel (1989), we hypothesize that the use of these markers by LMs will also have an impact on human-AI interactions.
@@
Our work begins with an examination of how LMs communicate uncertainties to end-users in realistic information seeking scenarios (§3). Specifically, we elicit responses from popular, publicly-deployed models including GPT, LLaMA-2, and Claude by prompting them to provide epistemic markers when answering multiple choice questions Brown et al. (2020); OpenAI (2023); Touvron et al. (2023); Anthropic (2022). Our analysis reveals that LMs are reluctant to share model uncertainties, despite errors in their generations. LMs can be explicitly prompted to use epistemic markers, but are more likely to generate expressions of certainty than uncertainty, despite an average 47% error rate among high confidence responses, (e.g., I’m confident the capital of Tanzania is Dar es Salaam. [Incorrect]).
We then investigate the behavioral responses of users towards model-generated epistemic markers (§4). While linguists and psychologists have long focused on the interpretation of epistemic markers by humans Budescu et al. (1988); Windschitl and Wells (1996); Wallsten et al. (1986), the pragmatic implications of the speakers of such markers being AI systems, combined with the known human over-reliance on AI Bussone et al. (2015); Jacobs et al. (2021); Bansal et al. (2021b); Buçinca et al. (2021), could drastically change their interpretation compared to human-spoken ones. Thus, we conduct several user studies to measure how individuals interpret and respond to uncertainties articulated by LMs in calibrated and miscalibrated settings (Figure 1). Our findings surprisingly indicate that users are heavily reliant on LM generated expressions of marked "I’m sure it’s…") and unmarked certainty (e.g., "The answer is…"). Subsequent experiments show that even minor miscalibrations in how a model uses epistemic markers can lead to long-term harms in human performance.
Lastly, given our findings on model overconfidence and human reliability of LM generations, we pinpoint the origins of model overconfidence (§5). We investigate model artifacts such as base models, instruction-tuned models, reward models, and human feedback datasets to isolate the origins of model confidence. Our investigation identifies the process of aligning models with human feedback (i.e., RLHF) as a key contributing factor and we uncover that human annotators are biased against expressions of uncertainty.
@@ 2 Epistemic Markers in Language Models
Linguists has extensively studied epistemic markers as ubiquitous linguistic features that signal speaker commitment and stance. These markers broadly fall into two categories: weakeners—expressions of uncertainty, and strengtheners—expressions of certainty Lakoff (1975); Hyland (2005, 2014).
In machine learning, work has focused on improving model calibration Jiang et al. (2021); Desai and Durrett (2020); Jagannatha and Yu (2020); Kamath et al. (2020); Kong et al. (2020) by calibrating the confidence value assigned by a model and model accuracy through a measure called ECE Naeini et al. (2015). Recent work has focused explicitly on how pretraining Hendrycks et al. (2019) and scaling Srivastava et al. (2022); Chen et al. (2023) impacts the calibration of language models. Most relevant to our work is Dhuliawala et al. (2023)’s studies on how humans interpret numerical confidences in calibrated and miscalibrated settings. A key issue remains, as numeric confidence values are known to be challenging for users to interpret Miller (2019). Our work, in contrast, aims for a more comprehensive understanding of how humans interpret LM-generated verbal epistemic markers.
We begin by eliciting open-ended generations, a departure from prior methods which prompts models to produce a predefined set of confidence expressions either numerically Kadavath et al. (2022); Tian et al. (2023); Liu et al. (2023); Xiong et al. (2024); Tanneru et al. (2023) or using an ordinal scale Lin et al. (2022); Mielke et al. (2020). Instead, our strategy enables a qualitative bottom-up Charmaz (2006) approach towards understanding how LMs generate epistemic markers, mimicking how real-world users might engage with LMs.
@@ 3.1 Methods
Our objective is to assess the potential harms and safety risks associated with widely used publicly deployed models like GPT, LLaMA-2, and Claude. [^1] Models were accessed June - November 2023. We pose a diverse set of questions from the Massive Multitask Language Understanding benchmark (MMLU Hendrycks et al., 2021), a four-way multiple-choice dataset spanning 57 subjects that assess both language model knowledge and problem-solving skills. We design confidence eliciting prompts and measure systematic trends in how LMs use strengtheners and weakeners.
@@
To ensure the generalizability of our results, we employ snowball sampling to generate a list of diverse prompts, gathering additional paraphrases of prompts from Amazon Mechanical Turk Workers and GPT-3.5 (details in §A).
We prompt nine models (text-davinci-003, GPT-3.5-Turbo, GPT-4, LLaMA-2 7B, LLaMA-2 13B, LLaMA-2 70B, Claude-1, Claude-2, Claude- Instant-1) using 49 prompts on 284 questions, resulting in a total of 125,244 queries. [^2] Duplicated question in the development set of MMLU benchmark, resulting in 284 instead of 285 questions. Using a zero-shot prompting approach, we aim to simulate the interactions of end-users. We set the temperature to 0.3 OpenAI (2023); Anil et al. (2023), with no stop tokens, and limit the token generation length to a maximum of 400 tokens (details in §C).
Eliciting and Classifying Epistemic Markers
The authors then qualitatively code Auerbach and Silverstein (2003); Charmaz (2006) for epistemic markers in generated responses from each model family via Regex pattern matching and categorizing them as strengtheners and weakeners. [^3] A future iteration could include training a classifier to identify codes at the trade-off of lowered transparency and interpretability.
Specifically, the process involved authors 1) manually looking through the generated responses, 2) identifying codes, 3) designing Regex heuristics to automatically detect markers, and 4) evaluating markers manually. Each evaluation consisted of randomly sampled 100 codes and evaluating them by hand. The entire qualitative coding process was repeated six times until we reached over 90% accuracy. Although this process involved the analysis of over a dozen models, as with most supervised methods, future researchers may need to adapt our schema when using it for out-of-distribution content. Our qualitative coding yielded a total of 76 strengtheners and 105 weakeners. See Tables 7 and 8 for the most commonly generated expressions.
3.2 Findings
@@
When using chain-of-thought instructions or explicit instructions to elicit epistemic markers, LMs can be encouraged to produce more epistemic markers. Resulting in 16% and 65% [^4] Weighted average emittance of epistemic markers when prompted with epistemic instructions and epistemic instructions with chain-of-thought of generations incorporating epistemic markers respectively.
@@
Six out of nine models have a preference to generate significantly more strengtheners than weakeners. Our results indicate that an average of 20% of generations had strengtheners while only 14% had weakeners. This bias is true among prompts eliciting for certainty, with or without CoT (Table 1). We see this trend emerge strongly among GPT and LLaMA-2 chat models, with the Claude-2 being more balanced in its generation of strengtheners and weakeners (Figure 3). Interestingly, the smaller models (LLaMA-2-7B and Claude-Instant-V1555Announced as a ”lighter” version of Claude. See: https://www.anthropic.com/index/introducing-claude) have a higher use of weakeners over strengtheners; model size and generation of epistemic markers could be explored in future work.
@@
Across all generations, only 53% of generations with expressions of certainty are correct (random accuracy being 25%). Although this accuracy rate is higher than accuracies among weakeners (32%), this is an alarmingly high rate of errors among strengtheners. Furthermore, due to the high rate of generations of strengtheners, 17% of all incorrect answers include strengtheners.
@@ 3.3 Discussion
The linguistic miscalibration is emerging as a new safety risk as LMs play a bigger role in human-LM collaborations. In the space of information seeking, the need for uncertainties and knowledge limitations will likely continue to persist even as model performance improves. Moving forward, we must examine how to mitigate the harms of model overconfidence and how to best build cognitive forcing designs, such as verbalized uncertainties, to discourage human overreliance of AI systems Buçinca et al. (2021).
@@ 4.1 Methods
For details on recruitment process, IRB protocol, and template selection see Appendix B.
@@ 4.2 Experimental Settings
Figure 4: Example of Setting 1 human experiments task.
Figure 5: Participant results in the calibrated, overconfident, and underconfident settings. We see lower scores across the miscalibrated rounds. In the overconfident setting, lowered scored persist in the later calibrated rounds as well.
@@
We first use our task to evaluate how participants rely on LM generated epistemic markers (from §3). Participants are shown a set of trivia questions and the beginning of a response (e.g., "I think it’s…" Figure 4). Users are asked whether or not they’d like to rely on Marvin’s answer. Since the users do not see any answers, they are simply expressing their reliance of epistemic markers as generated by LMs. Participants are presented with strengtheners, weakeners, and plain statements (expressions free of epistemic markers like, “The answer is (A)”). We recruited 25 participants and each were shown 106 questions.
@@
In these experimental settings, we introduce a scoring system, also modified from Bansal et al. 2019a. The scoring set-up is designed such that the only way to have a positive score is to rely on Marvin correctly based on relying the epistemic markers (see Table 2). [^8] To avoid spammers, we filtered the bottom 20% participants based on their performance on this task.
@@ 4.3 Findings
In the control setting, when presented with expressions of strengtheners, nearly 90% of users will rely on Marvin’s response. When presented with weakeners, approximately 90% of users will choose to look up the answer (see Table 6 for details). Surprisingly, plain statements like “The answer is (A)” or “(A)” are also relied on by users nearly 90% of the time. In other words, without communicating any epistemic markers, humans interpret this as a sign of model certainty.
@@
In the interactive calibrated setting, our results illustrate that users are able to learn mental models of epistemic markers after approximately 20 rounds. In the early rounds, users rely on strengtheners 94% of the time and on weakeners 7% of the time. After 20 rounds, most participants are able to nearly perfectly leverage Marvin’s epistemic markers with users learning to rely on strengtheners 99% of the time and weakeners nearly 1% of the time, averaging an accuracy of 97% across all rounds. The high performance in this calibrated setting also validates that participants are primarily relying on epistemic markers, to answer these questions. If they had been only relying on their own knowledge, it would be unlikely to see participants perform nearly perfectly on such challenging trivia questions.
@@
In the first 20 rounds, users relied on 81% of strengtheners, when only 66% of the generations with strengtheners were correct, accruing on greater penalties than necessary. Because participants are unlikely to know the answer to the question, we see that participants mistakenly rely on incorrect, confident generations an average 73% of the time. The consequences of system miscalibration continued to negatively impact human performance, even after several rounds of calibrated model answers. Participants averaged 76% in the miscalibrated rounds and 86% on the calibrated rounds (Figure 5).
@@
An unexpected effect of the overconfident responses was that participants started to interpret weakeners differently. Even though weakeners were used correctly in this setting (i.e., none of the answers with weakeners were correct), the participants were observed to rely on answers with weakeners at a higher rate than in the control setting (9% vs. 3%).
@@
Participants in the underconfident setting averaged 66% in the miscalibrated rounds but 98% on the later calibrated rounds, matching the performance in the calibrated setting. This is in contrast to the overconfident setting where users’ mental models were never fully corrected, even after the same number of calibrated rounds.